# Structured Generative Modeling with The Thermodynamic Kolmogorov-Arnold Model

### v5

@@ Abstract
Learning an energy-based model (EBM) in the latent space of a top-down generative model offers a versatile framework for generation across multiple data modalities. However, it remains unclear how its interpretability can be used to guide model design, improve generative quality, and reduce training time. Moreover, the reliance on Langevin Monte Carlo (LMC) sampling presents challenges in efficiency and sampling multimodal latent distributions. In this work, we propose a novel adaptation of the Kolmogorov-Arnold representation theorem for generative modeling and introduce the Thermodynamic Kolmogorov-Arnold Model (T-KAM) to take advantage of structural and inductive biases. By constraining the prior to univariate relationships, T-KAM enables fast and exact inference via the inverse transform method. With the low dimensionality of the latent space and suitable inductive biases encoded, we demonstrate that importance sampling (IS) becomes a viable, unbiased, and highly efficient posterior sampler. For situations where IS fails, we investigate a novel strategy using population-based LMC, which decomposes posterior sampling into a sequence of annealed distributions to improve multimodal sampling. T-KAM elegantly balances common trade-offs in generative modeling, offering fast inference, interpretability, and stable training, while being naturally suited to upcoming Zettascale Computing Corp. hardware.
@@ 1.1 Inductive biases and fast sampling
To derive our new model, which we call the Thermodynamic Kolmogorov-Arnold Model (T-KAM), we first introduce a novel adaptation of the Kolmogorov-Arnold Representation Theorem (KART) (Givental et al., 2009) by reinterpreting its inner functions as Markov kernels between probability spaces. Given its finite and definitive nature, KART can provide structural bias and reduce the reliance of architecture design on intuition and automated tuning.
@@ 1.2 Population-based Langevin sampling and Thermodynamic training
However, while ITS enables fast inference, posterior sampling must still be conducted using ULA when the IS estimator is insufficient. Prior work addressing poor MCMC convergence and multimodal exploration in EBMs has typically relied on diffusion models, such as the hierarchical EBM by Cui & Han (2024) and the diffusion-assisted EBM by Zhang et al. (2023). To maintain prior interpretability and fast inference post-training, we instead introduce posterior annealing as a novel and alternative strategy. This involves decomposing the posterior into a sequence of power posteriors, presented by Friel & Pettitt (2008), and recasting ULA with population-based MCMC methods (Swendsen & Wang, 1986; Marinari & Parisi, 1992; Hukushima & Nemoto, 1996).
The training criterion is derived from the discretized Thermodynamic Integral and the Steppingstone Estimator (SE), as presented by Calderhead & Girolami (2009); Annis et al. (2019) to ensure the impact of annealing on learning dynamics and parameter updates. Although this might be less expressive than hierarchical EBMs, it ensures that inductive biases remain easy to embed into the latent prior and interpret after training; and unlike denoising, annealing is embarrassingly parallel.
@@
MLE and SE are compared in Sec. 3.2. Despite our optimizations, the element-wise and localized nature of KANs complicate their parallelization on GPUs, which are best suited to matrix arithmetic. As a result, forward passes through T-KAM and the gradient evaluations required for ULA were found to be slow, necessitating reduced experimental complexity.
The experiments in Sec. 3.2 demonstrate that our mixture prior model is able to generate coherent, (although low-quality), samples after a few training epochs on RGB images, while still supporting efficient, exact sampling via component-wise ITS. In contrast, annealing incurred substantial computational overhead, but produced little to no improvements in sample quality, rendering its use difficult to justify. However, this result may change as a larger number of power posteriors, longer training durations, and more advanced posterior sampling algorithms become practical with the hardware advances proposed by Zetta (2024). Their architecture enables fine-grained control over data flow and parallelized element-wise nonlinearity which is naturally well suited to supporting T-KAM at scale, while improving the potential of univariate priors with its function fusion features.
We present T-KAM as an initial step to inspire further adaptations of KART in machine learning. This work aims to lay the groundwork for more exploration amongst the research community, with the far-reaching hope of eventually advancing towards a broader conceptualization: “The Kolmogorov-Arnold Representation Theorem Is All You Need.“
2 The Thermodynamic Kolmogorov-Arnold Model
@@
The Kolmogorov-Arnold Representation theorem (KART) (Givental et al., 2009) forms the basis of T-KAM. A structured proof of the theorem has been provided by Dzhenzher & Skopenkov (2022).
@@
Eq. 1 can provide structural bias and guide the design of T-KAM’s architecture. To ensure adherence to it, any transformations must be derived to be continuous, measurable, and deterministic. For readers less familiar with the technical background, the two most practically significant equations of Sec. 2.1 are Eqs. 57.
@@ 2.1 Univariate energy-based prior
We interpret $u_{p}\sim\mathcal{U}(u_{p};\,0,1)$ in Eq. 1 as a uniform random variable in the probability space, $\left([0,1],\mathcal{B}([0,1]),\mu\right)$, where $\mathcal{B}\left([0,1]\right)$ is the Borel $\sigma$-algebra on $[0,1]$ and $\mu$ is the Lebesgue measure on $[0,1]$. To ensure that the transformation $u_{p}\mapsto\psi_{q,p}(u_{p})$ is well defined in a probabilistic setting, we require it to be a Markov kernel. For all Borel measurable sets $A\subseteq\mathcal{Z}$, where $\mathcal{Z}\subset\mathbb{R}$ is introduced as the latent space of T-KAM, the map
@@
With an valid form for the normalized CDF, inverse transform sampling (ITS) can be used to draw samples from $\pi_{q,p}$ as deterministic transformations of uniform noise. This preserves the deterministic structure of KART while enabling flexible, learnable latent distributions. Inverse transform sampling remains effective even for functions that are not Riemann integrable, provided they are integrable in the Lebesgue sense, typical for the functions in KAN literature.
@@ 2.3.1 Generator
The generator, $\mathcal{G}$, can be considered a separate generator network allowing for more flexible realizations, such as convolutional layers when adherence to Eq. 1 is not required. Deepening this generator, as guided by Liu et al. (2024) for the Kolmogorov-Arnold Network (KAN), does not replace KART as the innermost foundation of T-KAM:
@@ 2.4 Maximum likelihood training
The conditional dependency on the collection, $\bm{f},\mathbf{\Phi}$, represents all learnable elements of T-KAM. For each sample of $\bm{x}^{(b)}$, the log-marginal likelihood in Eq. 16 can be separated out into two log-distributions by first introducing the latent vector, $\bm{z}$, otherwise absent upon marginalization over the latent distribution. This is shown under Sec. A.4.2, and the resulting expectation exhibits a lack of dependence on sampling, as proven under Sec. A.4.3:
@@ 2.4.1 Log-prior
Tractable normalization is available through Eq. 7, however it may be more efficient to discard normalization in the learning gradient. In Sec. A.4.4, we derive the contrastive divergence learning gradient for T-KAM, which faster to train and more typical of energy-based modeling:
@@ 2.4.3 Prior inverse transform sampling
In Eqs. 9 &  20, $z_{q,p}$ is drawn from $\pi_{q,p}$ with inverse transform sampling, (ITS) (Devroye, 1986). ITS can be used to sample from a component of the latent distribution, $\pi_{q,p}$, made possible given the unique characteristics of T-KAM. Specifically, the univariate, finite-grid structure of the constituent functions enables quadrature/interpolating integral approximations.
@@ 3 Experiments
Our experiments are designed to highlight the unique advantages of T-KAM and to compare the proposed strategies for scaling and training. In Sec. 3.1, we demonstrate the effects of introducing different inductive biases through the choice of reference distributions in Eq. 5, as well as through the KAN basis functions used for $f_{q,p}$ and $\Phi_{o,q}$ with comparisons between Radial Basis Functions (RBFs) (Li, 2024) and Fourier bases (FFTs) (Xu et al., 2024). We show that, when appropriate biases are encoded, importance sampling becomes a viable and highly efficient training strategy. We also demonstrate that the model remains effective when strictly adhering to KART with independent energy-based priors, thereby validating the use of KART as a source of structural inductive bias.
In Sec. 3.2, we evaluate the mixture EBM prior, sampled via component-wise ITS and paired with CNN-decoder architectures, which are smaller variants of those used by Pang et al. (2020). We additionally compare T-KAM trained using single-posterior ULA sampling with the MLE criterion against a population-based ULA sampling strategy using the SE criterion. These comparisons assess the practicality of using mixture distributions for scaling T-KAM and using annealing to improve training for high-dimensional RGB images.
@@ 3.1.1 MNIST & FMNIST
The MNIST (Deng, 2012) and FMNIST (Xiao et al., 2017) datasets are grayscale datasets that can be modeled with the first likelihood model from Eq. 21. The first collection of benchmarks presented in Sec. A.2 were obtained using the hyperparameters and architecture specified for this experiment in Tab. 4. T-KAM was trained for 10 epochs with 50,000 training examples, an importance sample size of $N_{s}=100$, and a batch size of $N_{x}=100$, (corresponding to 5,000 parameter updates). Prior sampling was conducted with ITS and posterior expectations were estimated with IS. The architecture was guided by KART and independent priors were used.
@@
T-KAM generated a diverse set of samples when a non-restrictive reference prior was used, (Gaussian or no reference), thus validating importance sampling and independent energy-based priors for simple datasets. Subjectively, the best MNIST samples arose with no reference prior and RBF bases, whereas the best FMNIST samples came from a Gaussian reference prior with RBF bases, potentially because FMNIST has smoother features than MNIST.
@@
In Sec. A.8, examples of trained univariate latent distributions have been plotted to demonstrate T-KAM’s interpretability. Learning can be understood by how T-KAM reshaped its latent representation relative to the initial reference prior. For instance, lognormal priors showed little change, indicating posterior collapse, and FFTs caused much sharper relationships than the smoother RBFs.
@@
where $k(\bm{z})$ is the 2D permeability field, $u(\bm{z})$ represents the 2D pressure field, and $f(\bm{z})$ is a forcing function evaluated at $\bm{z}$. Generating new flow-pressure samples is a challenging task in machine learning. One approach is to use input permeability samples while employing the Fourier Neural Operator (FNO) (Li et al., 2021), to learn the solution operator mapping between permeability and flow pressure. However, in this study, we use the purely generative task to illustrate how weak inductive biases can influence what T-KAM learns.
T-KAM was trained to generate new Darcy flow pressures using a batch size and importance sample size of $N_{x}=100,N_{s}=100$, and the first likelihood model in Eq. 21. Training spanned 1,000 epochs with 1,000 training examples, resulting in 10,000 parameter updates. All other hyperparameters are provided in Tab. 5. Prior sampling was conducted with ITS and posterior expectations were estimated with IS. The architecture was guided by KART and independent priors were used.
@@ 3.2 Langevin Monte Carlo and Steppingstone estimator
T-KAM, when adherent to KART, can be interpreted as a dense or spline-based network. However, for spatial data, convolutional neural networks (CNNs) (LeCun et al., 2015), are more suitable as the generator, $\mathcal{G}$. In addition to their computational efficiency through weight sharing, CNNs offer a key advantage over dense networks due to their spatial inductive bias: sequences of convolutional layers implement shift-equivariant transformations that are well-suited for images.
We evaluated T-KAM using a CNN generator against two RGB image datasets at different resolutions, SVHN ($32\times 32$) (Netzer et al., 2011) and CelebA ($64\times 64$) (Liu et al., 2015). The first likelihood model in Eq. 21 was adopted for both datasets. Due to limited computational resources during this study, (detailed in Sec. A.2), training was shortened to 4,000 parameter updates, (70,000 were used in the study by Pang et al. (2020)), and smaller versions of the CNN decoders used by Pang et al. (2020) were adopted as the generator architectures.
A single optimizer (Adam) was used to train both models, rather than using separate optimizers for the prior and generator. FFT bases were used for the prior energy functions, and full architectural details are provided under Sec. A.5.1. Posterior sampling was conducted with 20 steps of ULA, and no reference prior was used, $\pi_{0}=1$. T-KAM was trained for 20 epochs with a batch size of $N_{x}=100$ on 20,000 training examples. All remaining hyperparameters are in Tab. 6. Annealing was conducted with $N_{t}=10$ power posteriors and the $p$ scheduling outlined in Sec. A.3.1.
@@
$\overline{\text{FID}}_{\infty}=120.53$ $\overline{\text{FID}}_{\infty}=167.79$
$\overline{\text{KID}}_{\infty}=0.1103$ $\overline{\text{KID}}_{\infty}=0.1644$
@@
$\overline{\text{FID}}_{\infty}=174.69$ $\overline{\text{FID}}_{\infty}=173.78$
$\overline{\text{KID}}_{\infty}=0.1917$ $\overline{\text{KID}}_{\infty}=0.1851$
@@
Our proposed SE training strategy resulted in slightly better image metrics for CelebA, but significantly worsened them with SVHN. However, from our subjective perspective, the generated images appeared more well-defined, yet less diverse under SE than under MLE. Ultimately, we cannot justify the Thermodynamic training strategy due to its substantial computational overhead. With more available VRAM than in this study, it would be possible to parallelize a greater number of temperatures than $N_{t}=10$, which could potentially improve quality, as suggested by the results of Calderhead & Girolami (2009).

### v6

# Kolmogorov-Arnold Energy Models: Fast and Interpretable Generative Modeling

@@ Abstract
Learning an energy-based model (EBM) in the latent space of a top-down generative model offers a powerful framework for generation across many data modalities. However, it remains unclear how its interpretability can be used to guide model design, improve generative quality, and reduce training time. Moreover, the reliance on Langevin Monte Carlo (LMC) sampling presents challenges in efficiency and sampling multimodal latent distributions. We propose a novel adaptation of the Kolmogorov-Arnold representation theorem for generative modeling and introduce the Kolmogorov-Arnold Energy Model (KAEM) to take advantage of structural and inductive biases. By constraining the prior to univariate relationships, KAEM enables fast and exact inference via the inverse transform method. With the low dimensionality of the latent space and suitable inductive biases encoded, we demonstrate that importance sampling (IS) becomes a viable, unbiased, and highly efficient posterior sampler. For domains where IS fails, we introduce a strategy based on population-based LMC, decomposing the posterior into a sequence of annealed distributions to improve LMC mixing. KAEM balances common generative modeling trade-offs, offering fast inference, interpretability, and stable training, while being naturally suited to Zettascale Computing hardware.
@@ 1.1 Inductive biases and fast sampling
To derive our new model, which we call the Kolmogorov-Arnold Energy Model (KAEM), we first introduce a novel adaptation of the Kolmogorov-Arnold Representation Theorem (KART) (Givental et al., 2009) by reinterpreting its inner functions as Markov kernels between probability spaces. Given its finite and definitive nature, KART can provide structural bias and reduce the reliance of architecture design on intuition and automated tuning.
@@ 1.2 Population-based Langevin sampling and Thermodynamic training
However, while ITS enables fast inference, posterior sampling must still be conducted using ULA when the IS estimator is insufficient. Prior work addressing poor MCMC convergence and multimodal exploration in EBMs has typically relied on diffusion models, such as the hierarchical EBM by Cui & Han (2024) and the diffusion-assisted EBM by Zhang et al. (2023). To maintain prior interpretability and fast inference post-training, we instead introduce posterior annealing as a novel and alternative strategy. This involves decomposing the posterior into a sequence of power posteriors, presented by Friel & Pettitt (2008), and recasting ULA with population-based MCMC methods (Swendsen & Wang, 1986; Marinari & Parisi, 1992; Hukushima & Nemoto, 1996). can The training criterion is derived from the discretized Thermodynamic Integral and the Steppingstone Estimator (SE), as presented by Calderhead & Girolami (2009); Annis et al. (2019) to ensure the impact of annealing on learning dynamics and parameter updates. Although this might be less expressive than hierarchical EBMs, it ensures that inductive biases remain easy to embed into the latent prior and interpret after training; and unlike denoising, which scales sequentially, annealing can scale by adding more temperature in parallel.
@@
MLE and SE are compared in Sec. 3.2. Despite our optimizations, the element-wise and localized nature of KANs complicate their parallelization on GPUs, which are best suited to matrix arithmetic. As a result, forward passes through KAEM and the gradient evaluations required for ULA were found to be slow, necessitating reduced experimental complexity.
The experiments in Sec. 3.2 demonstrate that our mixture prior model is able to generate coherent, (although low-quality), samples after a few training epochs on RGB images, while still supporting efficient, exact sampling via component-wise ITS. In contrast, annealing incurred substantial computational overhead, but produced little to no improvements in sample quality, rendering its use difficult to justify. However, this result may change as a larger number of power posteriors, longer training durations, and more advanced posterior sampling algorithms become practical with the hardware advances proposed by Zetta (2024). Their architecture enables fine-grained control over data flow and parallelized element-wise nonlinearity which is naturally well suited to supporting KAEM at scale, while improving the potential of univariate priors with its function fusion features.
We present KAEM as an initial step to inspire further adaptations of KART in machine learning. This work aims to lay the groundwork for more exploration amongst the research community, with the far-reaching hope of eventually advancing towards a broader conceptualization: “The Kolmogorov-Arnold Representation Theorem Is All You Need.“
2 The Kolmogorov-Arnold Energy Model
@@
The Kolmogorov-Arnold Representation theorem (KART) (Givental et al., 2009) forms the basis of KAEM. A structured proof of the theorem has been provided by Dzhenzher & Skopenkov (2022).
@@
Eq. 1 can provide structural bias and guide the design of KAEM’s architecture. To ensure adherence to it, any transformations must be derived to be continuous, measurable, and deterministic. For readers less familiar with the technical background, the two most practically significant equations of Sec. 2.1 are Eqs. 57.
@@ 2.1 Univariate energy-based prior
We interpret $u_{p}\sim\mathcal{U}(u_{p};\,0,1)$ in Eq. 1 as a uniform random variable in the probability space, $\left([0,1],\mathcal{B}([0,1]),\mu\right)$, where $\mathcal{B}\left([0,1]\right)$ is the Borel $\sigma$-algebra on $[0,1]$ and $\mu$ is the Lebesgue measure on $[0,1]$. To ensure that the transformation $u_{p}\mapsto\psi_{q,p}(u_{p})$ is well defined in a probabilistic setting, we require it to be a Markov kernel. For all Borel measurable sets $A\subseteq\mathcal{Z}$, where $\mathcal{Z}\subset\mathbb{R}$ is introduced as the latent space of KAEM, the map
@@
With a valid form for the normalized CDF, inverse transform sampling (ITS) can be used to draw samples from $\pi_{q,p}$ as deterministic transformations of uniform noise. This preserves the deterministic structure of KART while enabling flexible, learnable latent distributions. Inverse transform sampling remains effective even for functions that are not Riemann integrable, provided they are integrable in the Lebesgue sense, typical for the functions in KAN literature.
@@ 2.3.1 Generator
The generator, $\mathcal{G}$, can be considered a separate generator network allowing for more flexible realizations, such as convolutional layers when adherence to Eq. 1 is not required. Deepening this generator, as guided by Liu et al. (2024) for the Kolmogorov-Arnold Network (KAN), does not replace KART as the innermost foundation of KAEM:
@@ 2.4 Maximum likelihood training
The conditional dependency on the collection, $\bm{f},\mathbf{\Phi}$, represents all learnable elements of KAEM. For each sample of $\bm{x}^{(b)}$, the log-marginal likelihood in Eq. 16 can be separated out into two log-distributions by first introducing the latent vector, $\bm{z}$, otherwise absent upon marginalization over the latent distribution. This is shown under Sec. A.4.2, and the resulting expectation exhibits a lack of dependence on sampling, as proven under Sec. A.4.3:
@@ 2.4.1 Log-prior
Tractable normalization is available through Eq. 7, however it may be more efficient to discard normalization in the learning gradient. In Sec. A.4.4, we derive the contrastive divergence learning gradient for KAEM, which faster to train and more typical of energy-based modeling:
@@ 2.4.3 Prior inverse transform sampling
In Eqs. 9 &  20, $z_{q,p}$ is drawn from $\pi_{q,p}$ with inverse transform sampling, (ITS) (Devroye, 1986). ITS can be used to sample from a component of the latent distribution, $\pi_{q,p}$, made possible given the unique characteristics of KAEM. Specifically, the univariate, finite-grid structure of the constituent functions enables quadrature/interpolating integral approximations.
@@ 3 Experiments
Our experiments are designed to highlight the unique advantages of KAEM and to compare the proposed strategies for scaling and training. In Sec. 3.1, we demonstrate the effects of introducing different inductive biases through the choice of reference distributions in Eq. 5, as well as through the KAN basis functions used for $f_{q,p}$ and $\Phi_{o,q}$ with comparisons between Radial Basis Functions (RBFs) (Li, 2024) and Fourier bases (FFTs) (Xu et al., 2024). We show that, when appropriate biases are encoded, importance sampling becomes a viable and highly efficient training strategy. We also demonstrate that the model remains effective when strictly adhering to KART with independent energy-based priors, thereby validating the use of KART as a source of structural inductive bias.
In Sec. 3.2, we evaluate the mixture EBM prior, sampled via component-wise ITS and paired with CNN-decoder architectures, which are smaller variants of those used by Pang et al. (2020). We additionally compare KAEM trained using single-posterior ULA sampling with the MLE criterion against a population-based ULA sampling strategy using the SE criterion. These comparisons assess the practicality of using mixture distributions for scaling KAEM and using annealing to improve training for high-dimensional RGB images.
@@ 3.1.1 MNIST & FMNIST
The MNIST (Deng, 2012) and FMNIST (Xiao et al., 2017) datasets are grayscale datasets that can be modeled with the first likelihood model from Eq. 21. The first collection of benchmarks presented in Sec. A.2 were obtained using the hyperparameters and architecture specified for this experiment in Tab. 4. KAEM was trained for 10 epochs with 50,000 training examples, an importance sample size of $N_{s}=100$, and a batch size of $N_{x}=100$, (corresponding to 5,000 parameter updates). Prior sampling was conducted with ITS and posterior expectations were estimated with IS. The architecture was guided by KART and independent priors were used.
@@
KAEM generated a diverse set of samples when a non-restrictive reference prior was used, (Gaussian or no reference), thus validating importance sampling and independent energy-based priors for simple datasets. Subjectively, the best MNIST samples arose with no reference prior and RBF bases, whereas the best FMNIST samples came from a Gaussian reference prior with RBF bases, potentially because FMNIST has smoother features than MNIST.
@@
In Sec. A.8, examples of trained univariate latent distributions have been plotted to demonstrate KAEM’s interpretability. Learning can be understood by how KAEM reshaped its latent representation relative to the initial reference prior. For instance, lognormal priors showed little change, indicating posterior collapse, and FFTs caused much sharper relationships than the smoother RBFs.
@@
where $k(\bm{z})$ is the 2D permeability field, $u(\bm{z})$ represents the 2D pressure field, and $f(\bm{z})$ is a forcing function evaluated at $\bm{z}$. Generating new flow-pressure samples is a challenging task in machine learning. One approach is to use input permeability samples while employing the Fourier Neural Operator (FNO) (Li et al., 2021), to learn the solution operator mapping between permeability and flow pressure. However, in this study, we use the purely generative task to illustrate how weak inductive biases can influence what KAEM learns.
KAEM was trained to generate new Darcy flow pressures using a batch size and importance sample size of $N_{x}=100,N_{s}=100$, and the first likelihood model in Eq. 21. Training spanned 1,000 epochs with 1,000 training examples, resulting in 10,000 parameter updates. All other hyperparameters are provided in Tab. 5. Prior sampling was conducted with ITS and posterior expectations were estimated with IS. The architecture was guided by KART and independent priors were used.
@@ 3.2 Langevin Monte Carlo and Steppingstone estimator
KAEM, when adherent to KART, can be interpreted as a dense or spline-based network. However, for spatial data, convolutional neural networks (CNNs) (LeCun et al., 2015), are more suitable as the generator, $\mathcal{G}$. In addition to their computational efficiency through weight sharing, CNNs offer a key advantage over dense networks due to their spatial inductive bias: sequences of convolutional layers implement shift-equivariant transformations that are well-suited for images.
We evaluated KAEM using a CNN generator against two RGB image datasets at different resolutions, SVHN ($32\times 32$) (Netzer et al., 2011) and CelebA ($64\times 64$) (Liu et al., 2015). The first likelihood model in Eq. 21 was adopted for both datasets. Due to limited computational resources during this study, (detailed in Sec. A.2), training was shortened to 8,000 parameter updates, (70,000 were used in the study by Pang et al. (2020)), and significantly smaller versions of the CNN decoders used by Pang et al. (2020) were adopted as the generator architectures.
A single optimizer (Adam) was used to train both models, rather than using separate optimizers for the prior and generator. FFT bases were used for the prior energy functions, and full architectural details are provided under Sec. A.5.1. Posterior sampling was conducted with 40 steps of ULA, and no reference prior was used, $\pi_{0}=1$. KAEM was trained for 40 epochs with a batch size of $N_{x}=100$ on 20,000 training examples. All remaining hyperparameters are in Tab. 6. Annealing was conducted with $N_{t}=5$ power posteriors and the $p$ scheduling outlined in Sec. A.3.1.
@@
$\overline{\text{FID}}_{\infty}=108.40$ $\overline{\text{FID}}_{\infty}=131.85$
$\overline{\text{KID}}_{\infty}=0.0955$ $\overline{\text{KID}}_{\infty}=0.1188$
@@
$\overline{\text{FID}}_{\infty}=122.18$ $\overline{\text{FID}}_{\infty}=127.54$
$\overline{\text{KID}}_{\infty}=0.1349$ $\overline{\text{KID}}_{\infty}=0.1420$
@@
Our proposed SE training strategy resulted in worse image metrics for both datasets. However, from our subjective perspective, the generated images appeared more well-defined, yet less diverse under SE than under MLE. Ultimately, we cannot justify the Thermodynamic training strategy due to its substantial computational overhead. With more VRAM than available in this study, it would be possible to parallelize a greater number of temperatures than $N_{t}=5$, which could potentially improve quality, as suggested by the results of Calderhead & Girolami (2009).